# SOC-DGL: Social Interaction Behavior Inspired Dual Graph Learning Framework for Drug-Target Interaction Identification

### v1

@@ 4.7 Hyperparameter Tuning Experiments
To assess the model’s robustness across multiple benchmark datasets, this study focused on the hyperparameters within the ADGL and EDGL modules and performed ”warm-start” experiments to systematically optimize these parameters.
@@ 4.7.1 Optimization of ADGL Hyperparameters
The number of convolutional layers ($N$) in the ADGL module was a critical hyperparameter influencing the model’s ability to capture the topological structure of drugs and targets. To assess the impact of different layer configurations on DTI prediction, experiments were conducted on both balanced and imbalanced versions of four benchmark datasets, with performance evaluated across multiple metrics, as shown in Fig. 5 (A-D). The results indicated that a two-layer GCN consistently achieved the best performance, particularly on the BindingDB, DrugBank, and Davis datasets. For $N=2$, the model excelled in AUROC, AUPR, and F1 score. However, on the KIBA dataset, $N=1$ slightly outperformed $N=2$ in AUROC (0.9838 vs. 0.9761) and AUPR (0.9872 vs. 0.9805), suggesting that a shallower GCN might have better captured global associations in this specific dataset. Overall, the two-layer setting expanded the receptive field, enhancing message passing and model expressiveness, while higher layer counts introduced complexity, increasing the risk of overfitting.
@@ 4.7.2 Optimization of EDGL Hyperparameters
This study focused on tuning key hyperparameters within the EDGL module. Specifically, this study investigated the total number of iterations ($K$) and the influence parameter ($\omega_{\alpha}$, denoted as $\alpha$) using the KIBA dataset. To thoroughly assess the impact of these hyperparameters, this study experimented with a range of values for $K$ and $\alpha$, evaluating their effects on the performance of SOC-DGL. This study explored $K$ values between 100 and 500 and adjusted $\alpha$ within the range of 0.05 to 0.25. The results presented in Fig. 5(E) showed that optimal performance was achieved when $K=200$ and $\alpha=0.20$, yielding an AUROC of 0.9761. This showed that a balanced and optimal choice of $K$ and $\alpha$ improved performance, mitigating the negative impact of extreme values.
@@ 4.7.3 Optimization of Adjustable Parameter $\varpi$
Traditional imbalanced loss functions addressed the negative sample imbalance by adjusting the influence of positive samples based on the ratio of negative to positive samples. However, this approach did not provide an optimal solution for handling the imbalance between positive and negative samples. This seemingly reasonable correction may have resulted in an excessively large influence of positive sample loss on the total loss, leading to artificial positive sample imbalance. Therefore, an adjustable parameter was required to adjust the correction for positive samples, ensuring that the imbalance function effectively handled negative sample imbalance. This study investigated the effect of the adjustable parameter $\varpi$ on imbalance processing through a hyperparameter experiment with $\varpi\in\{0.1,0.2,0.3,0.4,0.5,0.6\}$. Results in Fig. 5 (F) showed that $\varpi=0.2$ achieved the best performance, highlighting the necessity of introducing an adjustable parameter of appropriate size to adjust the correction for positive samples.
Figure 6: Case study analysis. (A) SOC-DGL predicts the top 9 drugs targeting ABL1 and demonstrates the associations among them. (B) Verification of ADGL-Edge and EDGL-Edge between Debromohymenialdisine and ABL1. (C) Perform a molecular docking experiment between Debromohymenialdisine and ABL1. The yellow dashed lines represent hydrogen bonds.
Table 3: SOC-DGL Prediction of the Top 10 Drugs Targeting Tyrosine-Protein Kinase ABL1.
@@
3 DB03878 N-[4-Methyl-3-[4-(3-Pyridinyl)-2-Pyrimidinyl]Amino]Phenyl]-3-Pyridinecarboxamide Berman et al. [34] 1 0
@@ 10 DB04367 Debromohymenialdisine None 3 1
4.8 Drugs prediction for the key target ABL1 in Chronic Myeloid Leukemi
To assess the potential of SOC-DGL in discovering DTI in real-world applications, this study conducted a case study on Chronic Myeloid Leukemia (CML), a myeloproliferative neoplasm with an incidence rate of approximately 1-2 cases per 100,000 adults, accounting for around 15% of newly diagnosed adult leukemia cases [38]. The primary pathogenesis of CML is driven by the fusion of the ABL1 gene on chromosome 9, a tyrosine kinase, with the Breakpoint Cluster Region (BCR) gene on chromosome 22, resulting in the formation of a malignant target that initiates and progresses the disease [39]. The ABL1 gene encodes a tyrosine kinase enzyme that activates downstream signaling pathways, regulating cell proliferation and differentiation [40]. CML onset is closely associated with ABL1 mutations, which lead to excessive cell proliferation, making it a key pathogenic factor in the disease [41].
[…]
Notably, SOC-DGL predicted Debromohymenialdisine as the 10th-ranked drug, which is a cyclin-dependent kinase inhibitor with pharmacological activity in the treatment of osteoarthritis and Alzheimer’s disease [42]. Upon searching databases and relevant literature, this study found that the direct association between Debromohymenialdisine and ABL1 had not been confirmed by any database or publication. However, SOC-DGL’s prediction suggested a potential link between the two. As shown in Fig. 6 (B) and Tab. 3, further investigation revealed that SOC-DGL first learned three direct similarity relations between DB04367, DB12267, and DB03878 through the ADGL module, as well as an interaction between DB04367 and Q07912. Subsequently, through the EDGL module, a friendship pathway was constructed using ABL1 as an intermediary, indirectly enhancing the higher-order similarity information between DB12267 and DB03878, and thereby uncovering a potential connection between ABL1 and DB04367. To further verify the binding degree of DB04367 and ABL1 at the molecular level, we conducted molecular docking experiments on them with Pymol and AutoDockTools, shown in Fig. 6 (C). The docking results indicated that DB04367 formed intermolecular forces with the amino acid residues in ABL1, further confirming the interaction between DB04367 and ABL1. These findings highlighted the remarkable ability of SOC-DGL to identify potential DTI through two distinct drug-target friendship strategies and validated the predictions of SOC-DGL, demonstrating its reliability in recognizing potential DTI and providing valuable insights for biological research.
@@ 5 Conclusion
In this study, SOC-DGL, a novel graph learning framework was proposed for DTI prediction. Drawing inspiration from real-world social interactions, this paper designed ADGL, which captures the direct global association and similarity information between drugs and targets based on extensive social connections. Additionally, EDGL leverages an even-polynomial graph filter driven by balance theory to extract higher-order similarity information from the heterogeneous drug-target graph, akin to high-order social interactions. This dual approach enhances prediction accuracy and demonstrates robust performance under data imbalance conditions. Experimental results on four benchmark datasets demonstrate that SOC-DGL outperforms existing state-of-the-art methods in both balanced and imbalanced data settings. The success of SOC-DGL in potential DTI (between DB01254 and ABL1) prediction highlights the importance of leveraging homogenous information aggregation in capturing complex indirect relationships within biological networks. Although SOC-DGL performs well in DTI prediction, it’s performance still limited by data quality, and future work could involve incorporating more biological knowledge or employing advanced data augmentation techniques to reduce noise and bias in datasets.
@@ 6.1.4 DrugBank dataset
6.1.5 Preprocessing of datasets
For each benchmark dataset, we extract various features of drugs and targets based on the SMILES sequences of drugs and the UniProt IDs of targets. For drug sequences, we use the RDKit software package to extract Molecular Access System (MACCS) keys fingerprints [45], topological (TOP) fingerprints [46], and Morgan fingerprints [47], which represent three distinct aspects of the molecular features. For target sequences, to provide a more systematic description of target sequences and their physicochemical properties, we employ the iLearn software package to calculate amino acid composition (AAC), Moran autocorrelation (MORAN) [48], Composition/Transition/Distribution (CTD) [49], and pseudo-amino acid composition (PAAC) [50], which encompass four different aspects of target features.
6.2 Detailed Process of Multi-view Learning
@@
Then, we define $v=v_{1}+v_{2}+v_{3}$ to represent the total number of perspectives. Following this, we introduce four auxiliary matrices $A$, $C1$, $C2$, and $C3$, each of which is a three-dimensional matrix of size $v\times n_{d}\times n_{d}$. These four auxiliary matrices encompass $v$ perspectives (distinct feature sets within multi-view learning), with each perspective having matrices of identical row and column dimensions. Specifically, $A$ denotes the relationship matrix among the various views, $C1$ is the auxiliary matrix for different views after singular value thresholding, $C2$ is the auxiliary matrix for different views following sparsity processing, and $C3$ is the auxiliary matrix that integrates information from multiple views. These auxiliary matrices are initialized to zero. Concurrently, we incorporate four Lagrange multiplier matrices $\lambda 1\in\mathbb{R}^{(v+1)n_{d}\times n_{d}}$, $\lambda 2\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, $\lambda 3\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and $\lambda 4\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and set five hyperparameters $\mu$, $\beta_{1}$, $\beta_{2}$, $\lambda$, and $\mu_{\text{max}}$, with only the parameter $\mu$ being continuously updated. For all formulas in this section, the subscript $i$ signifies the index of a particular view, where $i=0,1,\ldots,v$. We define $C_{\text{sum}}$, which represents the sum of the values of the $C2$ matrices of all views except view $i$ itself.
@@
LRSpNM 2021 0.9224 0.9382 0.8740 0.8740 0.8740 0.7730 0.8371
MLLMC 2022 0.9484 0.9592 0.9049 0.8935 0.9038 0.8803 0.9059
MULGA 2023 0.9479 0.9551 0.8984 0.8773 0.9404 0.7911 0.8600
MMDG-DTI 2025 0.9586 0.9646 0.9083 0.8949 0.9318 0.8483 0.8858
SOC-DGL \ 0.9761 0.9805 0.9350 0.9268 0.9373 0.9133 0.9328
LRSpNM 2021 0.9132 0.6904 0.6339 0.8998 0.6900 0.9299 0.5863
MLMC 2022 0.9329 0.7462 0.6874 0.9192 0.7084 0.9494 0.6675
MULGA 2023 0.8918 0.5556 0.5628 0.8666 0.6808 0.8934 0.4796
MMDG-DTI 2025 0.9245 0.7193 0.6639 0.9102 0.6998 0.9391 0.6273
SOC-DGL \ 0.9364 0.7473 0.6908 0.9207 0.7103 0.9506 0.6730
LRSpNM 2021 0.8930 0.9256 0.8644 0.8255 0.9083 0.6951 0.8250
MLMC 2022 0.9072 0.9395 0.8712 0.8712 0.9098 0.7177 0.8356
MULGA 2023 0.9158 0.9451 0.8790 0.8561 0.8539 0.8596 0.9058
MMDG-DTI 2025 0.9356 0.9647 0.8905 0.8627 0.9057 0.7934 0.8758
SOC-DGL \ 0.9467 0.9673 0.9038 0.8816 0.9097 0.8372 0.8985
LRSpNM 2021 0.9108 0.7067 0.6415 0.8961 0.6348 0.9409 0.6517
MLMC 2022 0.9110 0.7253 0.6704 0.9057 0.6530 0.9492 0.6888
MULGA 2023 0.8812 0.6179 0.5710 0.8698 0.5874 0.9187 0.5556
MMDG-DTI 2025 0.9087 0.7787 0.7369 0.9326 0.6503 0.9805 0.8409
SOC-DGL \ 0.9192 0.7655 0.6918 0.9134 0.6642 0.9561 0.7238
LRSpNM 2021 0.8514 0.5771 0.5533 0.8706 0.5464 0.9263 0.5602
MLMC 2022 0.8797 0.9285 0.8707 0.8750 0.8237 0.9286 0.9233
MULGA 2023 0.9874 0.9859 0.9678 0.9660 0.9818 0.9490 0.9542
MMDG-DTI 2025 0.9620 0.9661 0.9181 0.9172 0.9172 0.9340 0.9340
SOC-DGL \ 0.9894 0.9888 0.9641 0.9627 0.9677 0.9574 0.9606
LRSpNM 2021 0.8481 0.5727 0.5866 0.9096 0.6383 0.9399 0.5426
MLMC 2022 0.8992 0.8260 0.7990 0.9617 0.7584 0.9844 0.8443
MULGA 2023 0.9706 0.8366 0.7188 0.9444 0.6991 0.9721 0.7395
MMDG-DTI 2025 0.9149 0.7428 0.6967 0.8980 0.6504 0.9508 0.7429
SOC-DGL \ 0.9790 0.8705 0.8057 0.9594 0.8406 0.9726 0.7744
LRSpNM 2021 0.7563 0.8206 0.7016 0.6687 0.7522 0.5790 0.6573
MLMC 2022 0.8637 0.9126 0.8371 0.8415 0.7828 0.9052 0.8995
MULGA 2023 0.9715 0.9743 0.9165 0.9141 0.9112 0.9171 0.9219
MMDG-DTI 2025 0.9626 0.9658 0.9180 0.9159 0.9083 0.9240 0.9281
SOC-DGL \ 0.9735 0.9774 0.9286 0.9267 0.9208 0.9331 0.9369
LRSpNM 2021 0.7752 0.5480 0.5437 0.9257 0.4577 0.9758 0.6695
MLMC 2022 0.8594 0.7330 0.7043 0.9498 0.6181 0.9853 0.8185
MULGA 2023 0.9259 0.7179 0.6667 0.9358 0.6579 0.9659 0.6756
MMDG-DTI 2025 0.9051 0.6978 0.6717 0.9381 0.6522 0.9688 0.6943
SOC-DGL \ 0.9295 0.7654 0.7211 0.9508 0.6567 0.9824 0.8004
6.3 Comparison with Baseline Methods in “warm-start” experiments
To validate the effectiveness and superiority of SOC-DGL, we conducted comparative experiments with four leading methods that were widely recognized in DTI prediction. These models included GraphDTA [27], LRSpNM [28], MLMC [29], MULGA [30], and MMDG-DTI [31]. The detailed results of the comparative experiments on these baseline datasets were presented in Appendix Tab. 4.
@@ 6.3.2 LRSpNM
LRSpNM, introduced by Gaoyan Wu et al. in 2021, predicts drug-target interactions using similarity information and dynamic pre-filling of unknown interactions. It employs low-rank matrix approximation, loss functions, and Laplacian regularization, utilizing the ADMM algorithm for efficient computation on large bioinformatics datasets.
@@ 6.3.5 MMDG-DTI
6.4 Analysis of the even-polynomial graph filter in EDGL
Higher-order homogeneous information enhancement by the even-polynomial graph filter was primarily designed to capture direct relationships between adjacent nodes. In the context of a drug-target network, this approach emphasized identifying the direct interactions between drugs and targets. It effectively highlighted the interactions between first-order neighbors, which were nodes directly connected by an edge. Higher-order homogeneous information enhancement by the even-polynomial graph filter retained odd-order monomials while discarding even-order terms, as formulated below:
$Y^{\prime}=f(\hat{X},P)=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k-1}\hat{X}$ (25)
Regarding the self-attention mechanism, its application in graph neural networks captured the complex interactions between nodes in a graph. By calculating attention weights, the self-attention mechanism assigned different weights to different neighboring nodes, thereby selectively aggregating important neighboring information. It focused on the most critical relationships for the prediction task, whether direct or indirect. For the self-attention mechanism, we used the result $H^{\prime}$ obtained from ADGL as the input and performed three different linear transformations on the input vector to obtain Q (Query), K (Key), and V (Value), where $Q=H^{\prime}W^{Q},K=H^{\prime}W^{K},V=H^{\prime}W^{V}$. Here, $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d_{k}}$ were trainable weight matrices, and here $d_{k}=m$. Then, the formula for self-attention was obtained as:
$Y^{\prime\prime}=\text{softmax}(\frac{QK^{T}}{\sqrt{m}})\cdot V$ (26)

### v2

@@ 4.5 Hyperparameter Tuning Experiments
To assess the model’s robustness across multiple benchmark datasets, this study focused on three kinds of hyperparameters in SOC-DGL and systematically optimized these parameters.
@@ 4.5.1 Optimization of ADGL Hyperparameters
The number of convolutional layers ($N$) in the ADGL module significantly influences the ability to capture the topological structure information. To evaluate the impact of layer configurations on DTI prediction, experiments were conducted on both balanced and imbalanced versions of four benchmark datasets, as shown in Fig. 4 (A-D). Results demonstrated that a two-layer GCN consistently yielded the best performance. However, on the KIBA dataset, $N=1$ slightly outperformed $N=2$ in AUROC and AUPR, indicating that a shallower GCN better captured global associations in this specific case. In general, the two-layer configuration expanded the receptive field, enhancing message passing and model expressiveness, while additional layers increased complexity, raising the risk of overfitting.
@@ 4.5.2 Optimization of EDGL Hyperparameters
This study focused on optimizing key hyperparameters within the EDGL module, specifically the total number of iterations ($K$) and the influence parameter ($\alpha$), using the KIBA dataset. A range of values for $K$ (100 to 500) and $\alpha$ (0.05 to 0.25) were tested to evaluate their impact on SOC-DGL performance. As shown in Fig. 4 (E), optimal performance was achieved with $K=200$ and $\alpha=0.20$, resulting in an AUROC of 0.9761. These results indicate that carefully balancing $K$ and $\alpha$ enhances model performance, while extreme values adversely affect results.
@@ 4.5.3 Optimization of Adjustable Parameter $\varpi$
Traditional imbalanced loss functions address the imbalance by adjusting the influence of positive samples based on their ratio to negative samples. However, it can lead to an excessive influence of positive sample loss, creating artificial positive sample imbalance. To mitigate this, an adjustable parameter $\varpi$ was introduced to fine-tune the correction for positive samples. Hyperparameter experiments with $\varpi\in\{0.1,0.2,0.3,0.4,0.5,0.6\}$ were conducted, and the results in Fig. 4 (F) demonstrated that $\varpi=0.2$ achieved optimal performance. These findings emphasize the importance of appropriately tuning the adjustable parameter to effectively handle sample imbalance.
Figure 5: Case study analysis. (A) SOC-DGL predicts the top 9 drugs targeting ABL1 and demonstrates the associations among them. (B) Verification of ADGL-Edge and EDGL-Edge between DB04367 and ABL1. (C) Perform a molecular docking experiment between DB04367 and ABL1.
4.6 Case study for the ABL1 in Chronic Myeloid Leukemi
To assess the potential of SOC-DGL in discovering DTI in real-world applications, this study conducted a case study on Chronic Myeloid Leukemia (CML), a myeloproliferative neoplasm accounting for around 15% of newly diagnosed adult leukemia cases [41]. The primary pathogenesis of CML is driven by the fusion of the ABL1 gene on chromosome 9, a tyrosine kinase, with the Breakpoint Cluster Region (BCR) gene on chromosome 22, resulting in the formation of a malignant target that initiates and progresses the disease [42]. The onset of CML is closely associated with ABL1 mutations, which lead to excessive cell proliferation, making it a key pathogenic factor in the disease [43].
[…]
In particular, the 10th drug DB04367 is a cyclin-dependent kinase inhibitor with pharmacological activity in the treatment of osteoarthritis and Alzheimer’s disease [44]. However, no database or publication confirmed the direct association between DB04367 and ABL1. In this study, SOC-DGL suggested a possible link between them. As shown in Fig. 5 (B) and Table. 4, further investigation revealed that SOC-DGL first learned three direct similarity relations between DB04367, DB12267, and DB03878 through the ADGL module, as well as an interaction between DB04367 and Q07912. Subsequently, through the EDGL module, a friendship pathway was constructed using ABL1 as an intermediary, indirectly improving the higher-order similarity information between DB12267 and DB03878, and thus uncovering a potential connection between ABL1 and DB04367.
To verify the affinity of DB04367 and ABL1 at the molecular level, we conducted molecular coupling experiments on them with Pymol and AutoDockTools, shown in Fig. 5 (C). The results indicated that DB04367 formed intermolecular forces with an intermolecular hydrogen bond in ABL1, confirming the interaction between DB04367 and ABL1.
Table 4: SOC-DGL Prediction of the Top 10 Drugs Targeting ABL1.
@@ 4.6 Case study for the ABL1 in Chronic Myeloid Leukemi
3 DB03878 NP-3-PP(a) Berman et al. [47] 1 0
@@ 10 DB04367 Debromohymenialdisine None 3 1
  • (a)
    Full name: N-[4-Methyl-3-[4-(3-Pyridinyl)-2-Pyrimidinyl]Amino]Phenyl]-3-Pyridinecarboxamide.
Table 5: Perform Similarity Ensemble Approach on the drug DB04367.
Target Key Target Name Description P-Value $\downarrow$ MaxTc $\uparrow$
CHK2_HUMAN CHEK2 Serine/threonine-protein kinase Chk2 1.422e-09 0.62
CHK1_HUMAN CHEK1 Serine/threonine-protein kinase Chk1 1.826e-09 0.62
UBP7_HUMAN USP7 Ubiquitin carboxyl-terminal hydrolase 7 1.315e-07 0.63
To further verify the potential association between DB04367 and ABL1 from the perspective of biological pathways, we used the SEA (Similarity Ensemble Approach) [51] to predict the potential targets of DB04367. SEA is a method based on chemical structure similarity, which infers possible targets by comparing the chemical structure of the query compound with known compounds in curated databases. Results shown in Table 5 suggest that DB04367 may interact with CHEK2, CHEK1, and USP7. These P-values indicate strong associations between the compound and known ligands of the predicted targets. Related papers indicate that the above predicted targets may be biologically associated with ABL1: (1) CHEK2 and ABL1 are potentially linked in the DNA damage response pathway, particularly through coordinated regulation in EWS-mediated alternative splicing [52]; (2) CHEK1 and ABL1 are co-enriched in multiple Gene Ontology biological processes related to RAD51, such as homologous recombination repair of DNA double-strand breaks and regulation of cell proliferation [53]; and (3) USP7 has been shown to enhance the stability of the BCR-ABL1 fusion protein, thereby promoting the proliferation and drug resistance of CML cells [54]. These findings provide biological evidence for the potential association between DB04367 and ABL1, supporting the reliability of SOC-DGL’s prediction.
@@ 5 Conclusion
This study proposes SOC-DGL, a novel graph learning framework for DTI prediction. Inspired by real-world social interactions, SOC-DGL introduces ADGL to capture direct global associations and similarities between drugs and targets through extensive social connections. Additionally, EDGL employs a balance theory-driven even-polynomial graph filter to extract higher-order similarities from the heterogeneous drug-target graph, reflecting complex, higher-order interactions. Experimental results on four benchmark datasets demonstrate that SOC-DGL outperforms SOTA methods under both balanced and imbalanced conditions. Notably, SOC-DGL’s successful prediction of the DTI between DB01254 and ABL1 emphasizes the importance of homogeneous information aggregation in uncovering complex indirect relationships in biological networks. We anticipate that future work will incorporate additional biological knowledge or advanced data augmentation techniques to address dataset noise and bias.
@@ 6.1.4 DrugBank dataset
6.2 Methodological Supplement
6.2.1 Detailed Process of Multi-view Learning
@@
Then, we define $v=v_{1}+v_{2}+v_{3}$ to represent the total number of perspectives. Following this, we introduce four auxiliary matrices $A$, $C1$, $C2$, and $C3$, each of which is a three-dimensional matrix of size $v\times n_{d}\times n_{d}$. These four auxiliary matrices encompass $v$ perspectives (distinct feature sets within multi-view learning), with each perspective having matrices of identical row and column dimensions. Specifically, $A$ denotes the relationship matrix among the various views, $C1$ is the auxiliary matrix for different views after singular value thresholding, $C2$ is the auxiliary matrix for different views following sparsity processing, and $C3$ is the auxiliary matrix that integrates information from multiple views. These auxiliary matrices are initialized to zero. Concurrently, we incorporate four Lagrange multiplier matrices $\lambda 1\in\mathbb{R}^{(v+1)\times n_{d}\times n_{d}}$, $\lambda 2\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, $\lambda 3\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and $\lambda 4\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and set five hyperparameters $\mu$, $\beta_{1}$, $\beta_{2}$, $\lambda$, and $\mu_{\text{max}}$, with only the parameter $\mu$ being continuously updated. For all formulas in this section, the subscript $i$ signifies the index of a particular view, where $i=0,1,\ldots,v$. We define $C_{\text{sum}}$, which represents the sum of the values of the $C2$ matrices of all views except view $i$ itself.
@@
6.2.2 Residual Connection Techniques
The corresponding formula is as follows:
$\hat{H}=\omega H^{\prime}+(1-\omega)H^{\prime\prime}$ (21)
6.3 The details of Baseline Methods
To validate the effectiveness and superiority of SOC-DGL, we conducted comparative experiments with seven leading methods that were widely recognized in DTI prediction. These models included GraphDTA, LRSpNM, MLMC, MULGA, MSI-DTI, BCMMDA and MMDG-DTI.
@@ 6.3.2 LRSpNM
LRSpNM, introduced by Gaoyan Wu et al. in 2021, predicts DTI using similarity information and dynamic pre-filling of unknown interactions. It employs low-rank matrix approximation, loss functions, and Laplacian regularization, utilizing the ADMM algorithm for efficient computation on large bioinformatics datasets.
@@ 6.3.4 MULGA
6.3.5 MSI-DTI
MSI-DTI, proposed by Zhao et al. in 2024, utilizes multi-source information for DTI prediction. By integrating multi-head self-attention within a transformer architecture, it effectively captures complex interactions across diverse drug and target data sources, improving prediction accuracy and robustness.
6.3.6 BCMMDA
BCMMDA, introduced by Huang et al. in 2024, enhances drug–protein interaction prediction by leveraging branch-chain mining (BCM) for fragment extraction and a multi-dimensional attention mechanism. This method improves feature representation and prediction accuracy through deep convolutional neural networks (CNNs) and attention-based refinement.
@@ 6.3.7 MMDG-DTI
6.4 Metrics
To comprehensively assess SOC-DGL’s capability in handling drug-target repurposing, we set up five metrics: ACC, Precision, Recall, Specificity, and F1_score.
$\text{ACC}=\frac{TP+TN}{TP+TN+FN+FP}$ (22)
$\text{Precision}=\frac{TP}{TP+FP}$ (23)
$\text{Recall}=\frac{TP}{TP+FN}$ (24)
$\text{Specificity}=\frac{TN}{TN+FP}$ (25)
[…]
where TP, FP, FN, and TN represent the number of samples that are actually positive and predicted as positive (True Positive Samples), actually negative and predicted as positive (False Positive Samples), actually positive and predicted as negative (False Negative Samples), and actually negative and predicted as negative (True Negative Samples), respectively. Here, ”positive” indicates that there is an association between a drug and a target, while ”negative” indicates that there is no association (or very low association) between a drug and a target. Subsequently, based on the aforementioned metrics, we plot the ROC curve and the PR curve, and then calculate the area under the ROC curve (AUROC) and the area under the PR curve (AUPR). The five metrics mentioned earlier, along with the two curve area metrics, provide a total of seven indicators that can measure the model’s performance from different aspects, assisting in the selection and optimization of classification models to meet the needs of DTI prediction.